OpenStudy (jkbo):

Please help What is the slope of the line through the points (–2, –1) and (8, –3)?

jimthompson5910 (jim_thompson5910):

use the slope formula m = (y2 - y1)/(x2 - x1)

OpenStudy (a_clan):

You have (x1,y1) and (x2,y2) The standard formula for slope, then, is : m(slope) = (y2-y1)/(x2-x1) Which you can find in your book.

OpenStudy (anonymous):

Slope = (-3+1)/8+2= -2/10= -1/5